Your coin collection contains 40 1955 silver dollars. If your grandparents purchased them for their face value when they were new, how much will your collection be worth when you retire in 2059, assuming they appreciate at a 11 percent annual rate?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Hence

A=40(1.11)^104

=40*51711.81652

=2,068,472.66(Approx).